what is the height of a triangle if the base is 10 cm and both sides are 13 cm
Answer by amarjeeth123(575)


In the given triangle AB=AC=13 cm
Base BC=10 cm
BD=DC=BC/2=5 cm
Using Pythagoras theorem,
AD=√AC^2-DC^2=√13^2-5^2=√169-25=√144=12 cm
The height of the triangle is 12 cm.